OpenStudy (anonymous):

Find the LCD that would eliminate the fraction in this equation. 1/2y-2/3=7/8

OpenStudy (anonymous):

1/2y=7/8+2/3 1/2y=21+16/24 1/2y=37/24 y=37/12

OpenStudy (anonymous):

so is 37/12 what i have to multiply to both sides?

OpenStudy (anonymous):

wait?

OpenStudy (anonymous):

multiplay by 12

OpenStudy (anonymous):

and tell me the ans as well.thnks

OpenStudy (anonymous):

How can i tell you the answer if im the one who asked it

OpenStudy (anonymous):

Can someone just help me besides deadman340 because honestly the person keeps giving me the wrong answer

OpenStudy (anonymous):

dear 12 is the right ans.

OpenStudy (anonymous):

no its not, thats what you divide or multiply by the fractions

OpenStudy (agentnao):

The LCD will be 24y